# TrinityDNA: A Bio-Inspired Foundational Model for Efficient Long-Sequence DNA Modeling

### v1

@@ 1 Introduction
The rapid advancements in large-scale, long-sequence modeling, particularly in the realm of Natural Language Processing (NLP) (team2023gemini, ; achiam2023gpt, ), have radically transformed the way we approach complex data. Deep learning models, such as Transformers (vaswani2017attention, ), have achieved unprecedented success in tasks that span from language translation to text generation, revolutionizing not just NLP but a variety of other fields. These models have proven their capability to capture intricate dependencies in data, providing solutions to challenges that were once considered insurmountable. With these breakthroughs in NLP, there has emerged an exciting opportunity to extend the power of sequence modeling to a completely different domain—genomics—where data shares some key similarities, such as its sequential nature (zhou2015predicting, ).
Genomic data, particularly DNA sequences, consists of extraordinarily long strings of information that encode the fundamental building blocks of life. Unlike the highly dense and structured data typically encountered in NLP, genomic sequences are sparse in nature, containing vast stretches of repetition and variability (liu2024genbench, ). Despite this, they hold a rich repository of biological information that is crucial for understanding gene functions, regulatory mechanisms, and cellular processes. The ability to model DNA sequences deeply could lead to breakthrough applications in personalized medicine, genetic engineering, and the overall understanding of biological systems. However, effectively capturing the dependencies within such long, sparse sequences remains a significant challenge.
Figure 1: Overview of TrinityDNA Model: (Left) The evolutionary training strategy of TrinityDNA, progressing from prokaryotic DNA to multi-species eukaryotic DNA, and its DNA-targeted long-sequence modeling approach addressing structural features such as bidirectional complementarity and major/minor grooves. (Right) Radar chart illustrating the state-of-the-art performance on the zero-shot performance of our models versus popular models such as EVO and Caduceus. The TrinityMicroDNA model refers to a model trained only on prokaryotic data, which typically performs better on prokaryotic tasks, while TrinityDNA is better on eukaryotic tasks.
While the parallels between NLP and genomics are evident, directly applying traditional NLP models to genomic sequences proves difficult. The sparse, low-density nature of DNA sequences means that existing models often struggle to identify long-range dependencies and interpret the underlying biological structures(mallet2021reverse, ; zhou2021towards, ). Moreover, the lack of biologically informed features in current models limits their effectiveness in genomic contexts. Many models trained on single-species data perform poorly when generalized to other species or broader biological contexts. As a result, the impact of these models on genomic research has been somewhat restricted, and their applicability to real-world challenges remains limited.
@@
In summary, TrinityDNA represents a significant advancement in DNA sequence modeling, addressing critical limitations of existing models through innovative architectural designs, incorporation of biological insights, and robust training strategies. By bridging the gap between NLP advancements and genomic data intricacies, TrinityDNA paves the way for more accurate predictions of gene functions, enhanced understanding of genetic regulation mechanisms, and broader applications in genomics research.
@@ 2 Background
Deoxyribonucleic acid (DNA) is the hereditary material in most living organisms, consisting of long sequences of nucleotides. Each nucleotide comprises a phosphate group, a deoxyribose sugar, and one of four nitrogenous bases: adenine (A), thymine (T), cytosine (C), and guanine (G). A DNA sequence can be represented as a string $S=(s_{1},s_{2},\ldots,s_{N})$, where $s_{i}\in\{A,T,C,G\}$ and $N$ denotes the sequence length. Genomic sequences are characterized by their low-density encoding and considerable length, often extending to millions of nucleotides, which presents unique challenges for sequence modeling.
@@
The double helix structure of DNA features two distinct grooves: the minor groove and the major groove. These grooves arise from the asymmetric positioning of the phosphate backbone relative to the base pairs. (1) Major Groove: wider and deeper, the major groove provides greater accessibility for protein binding and molecular interactions. It generally covers five to seven nucleotides. (2) Minor Groove: narrower and shallower, the minor groove presents a different arrangement of hydrogen bond donors and acceptors. While less accessible than the major groove, its length is three to five nucleotides. Understanding the geometric properties of these grooves is essential for predicting functional regions within the genome.
@@
DNA molecules consist of two complementary strands running in opposite directions, a feature known as antiparallel orientation. For a DNA sequence $S=(s_{1},s_{2},\ldots,s_{N})$, its reverse complement $S^{R}$ is defined as:
@@
2.2 DNA Long-Sequence Modeling
@@
A prominent class of models for handling long-range dependencies is based on Structured State Space Models (SSMs) (gu2021efficiently, ; gu2021combining, ; gu2022parameterization, ; gupta2022diagonal, ; smith2022simplified, ; dao2022hungry, ). These models emerge from discretizing a continuous-time linear system:
@@
HyenaDNA (poli2023hyena, ) is an SSM variant, a decoder-style model capable of handling long genomic sequences (e.g., hundreds of thousands of tokens). It leverages the Hyena operator, which replaces traditional attention with a fast convolution-based mechanism. Concretely, HyenaDNA blocks compute a Toeplitz convolution filter on projected input tokens, allowing processing of very large contexts (in $\mathcal{O}(L\log L)$ time) without the quadratic cost typically associated with attention.
@@
MambaDNA (schiff2024caduceus, )—also referred to as Caduceus—builds on the selective SSM approach of Mamba and incorporates reverse-complement symmetry, a core property of DNA sequences. In standard Mamba, the module processes sequences in a single direction. MambaDNA extends this design in two ways: (1) BiMamba: Instead of only left-to-right processing, BiMamba applies the Mamba block twice. (2) RC Equivariance: MambaDNA explicitly enforces RC symmetry by taking a sequence and its RC as inputs to the same SSM-based module. See more details in Appendix D.
@@ 3 TrinityDNA: Sequence, Structure, and Strategy for DNA Modeling
[…]
3.1 Preliminaries
@@
While SSMs theoretically excel at handling long sequences, they inherently exhibit a locality bias (wang2024understand, ). This issue is amplified in DNA sequence modeling, where dependencies across vast stretches of genetic material need to be captured for accurate biological interpretation. Specifically, in genomic data, long-range dependencies often span tens of thousands or even hundreds of thousands of base pairs. Existing models like SSMs typically focus on local dependencies due to the computational challenges of processing long sequences. Consequently, our empirical results in Figure 2 demonstrate that the SSM-based model (Caduceus) lost their focus as sequence length increased, while the full-attention-based model (DNABERT2) suffered from heavy computation. These limitations hamper the ability of modern DNA foundation models to fully capture the complexities inherent in DNA sequences, particularly when dealing with biological functions and regulatory mechanisms that depend on large, interconnected genomic regions.
@@
As sequences grow in training, full self-attention “flattens out”: attention scores converge toward a uniform, high-entropy distribution in Figure 3, so every token is weighted almost equally, and useful signals vanish. This oversmoothing hits low-density data—images, DNA-hardest, where informative tokens are sparse and far apart. In shorter windows, the same model still shows specialized heads, e.g., retrieval vs. induction (xiao2024duo, ), but that diversity collapses at the kilobase scale. These findings motivate a multi-window, multi-head scheme that mixes narrow local windows with broader global ones to curb entropy and retain head specialization.
@@
3.2 Groove Fusion Module
@@
3.3 Sliding Multi-Window Attention
To overcome the locality bias and attention oversmoothing, we revisit the design of Multi-Head Attention (MHA), focusing on varying attention window sizes for different heads. In traditional MHA, each head typically attends to the entire sequence using a fixed attention window, which can limit the model’s ability to capture dependencies at different scales. In the context of DNA sequence modeling, dependencies are not uniform across the sequence. For example, short motifs require local attention, while regulatory regions may span longer stretches of DNA. To address this, we introduce a multi-scale attention grouping strategy named SMWA. In this approach, we assign different attention heads to focus on different scales of dependencies in the DNA sequence, allowing the model to specialize in either local or global dependencies depending on the scale of the feature being modeled. Formally, we define the window sizes for each attention head $h$ as $L_{h}\in\mathbb{N}$, with $L_{h}$ representing the length of the attention window for head $h$. For head $h$, the attention mechanism is computed using a sliding window of size $L_{h}$ over the sequence, allowing the model to focus on different scales. Specifically, the attention calculation is constrained to the sliding window around each position, which captures dependencies within a local region of the sequence. The sliding window can be defined as:
$\text{Attn}_{h}(S_{i})=\text{Softmax}\left(\frac{Q_{h}(i)K_{h}(i+[-L_{h},L_{h}])^{T}}{\sqrt{d_{k}}}\right)V_{h}(i+[-L_{h},L_{h}])$
where $Q_{h}\in\mathbb{R}^{N\times d_{k}}$ is the query matrix for head $h$, $K_{h}\in\mathbb{R}^{N\times d_{k}}$ is the key matrix, $V_{h}\in\mathbb{R}^{N\times d_{v}}$ is the value matrix, $N$ is the sequence length, and $d_{k}$ and $d_{v}$ are the dimensionality of the key and value vectors, respectively. $i$ is the index of the sequence, and $[-L_{h},L_{h}]$ represents the range of indices for the sliding window around $i$. This enables each head to specialize in attending to either short-range or long-range dependencies by adjusting $L_{h}$. The final output of the multi-head attention layer is the concatenation of all heads’ outputs, followed by a linear transformation:
@@ 3.3 Sliding Multi-Window Attention
3.4 Gated Reverse Complement
To leverage the reverse complementarity inherent in DNA sequences, we introduce a novel Gated Reverse Complement (GRC) mechanism. This mechanism employs a shared Transformer module to process both the forward and reverse complement sequences in parallel. The reverse complement of a DNA sequence $S=(s_{1},s_{2},\dots,s_{N})$ is defined as $S^{R}=(s_{N}^{C},s_{N-1}^{C},\dots,s_{1}^{C})$, where $s_{i}^{C}$ denotes the complementary base of $s_{i}$ (with base-pairing rules $A\leftrightarrow T$, $C\leftrightarrow G$). The GRC mechanism works by feeding both the forward and reverse complement sequences into a shared SMWA-equipped Transformer network $f_{\theta}$, where the outputs are gated using a linear gating mechanism to combine the two representations effectively. The gating at the final layer is given by:
@@ 3.4 Gated Reverse Complement
where $W_{G}$ are learned weights, $\sigma$ represents the sigmoid or identity function, and the Flip operator means to reverse the sequence as the original order. This allows the model to simultaneously learn the representations of both forward and reverse sequences, capitalizing on the symmetry of DNA, which is crucial for many biological processes.
3.5 Evolutionary Training Strategy
The Evolutionary Training Strategy (ETS) approach leverages a two-stage, evolution-inspired training strategy to progressively address the varying complexities of genomic data. In the first stage, the model is trained on prokaryotic genomes, which average around 924 base pairs (xu2006genome, ; nguyen2024sequence, ) and present relatively straightforward regulatory architectures. Through this foundational phase, the model captures essential DNA sequence motifs and organizational patterns. Subsequently, the second stage introduces eukaryotic genomes, known for their intron-exon structures and gene lengths that can span tens of kilobases (dalla2024nucleotide, ). Alongside this transition, the model’s context window is enlarged from 8K to 100K base pairs, accommodating multiple co-expressed genes and complex regulatory elements.
@@ 4.1 Pre-training
We adopt masked language modeling (MLM) (devlin2018bert, ) and character-level tokenization for all our DNA models. Following our Evolutionary training strategy, we conduct pre-training in two phases: Stage 1 (Short-Sequence Pre-training): We use prokaryotic (bacterial and archaeal) DNA data from the OpenGenome dataset introduced in nguyen2024sequence , training on sequences of length 8k to learn fundamental nucleic acid patterns in shorter contexts. Stage 2 (Multi-species Post-training): We then continue pre-training on a multi-species collection, the Multispecies dataset, presented in dalla2024nucleotide . The sequences in this stage can be as long as 100k, spanning archaebacteria, fungi, vertebrate genomes, and more. This step exposes the model to a rich spectrum of evolutionary signals, enabling it to handle much longer contexts and to capture the diverse structural intricacies of eukaryotic DNA. Hence, we propose two main models, TrinityMicroDNA and TrinityDNA, each with 1 billion (1B) parameters. The former is trained solely on prokaryotic data, while the latter builds upon the former by post-training on eukaryotic data. A “100k” label indicates an increased context window size of up to 100k. By transitioning from bacterial genomes to longer eukaryotic genomes, our method achieves broad coverage of genomic features across scales. See more details in Appendix A.
@@
4.2 Ablation Study and Analysis
Table 1: Effect of GRC, GFM, and SMWA on pre-training perplexity.
Components W/O W
GRC 2.731 2.599 (-0.132)
GFM 2.599 2.534 (-0.065)
SMWA 2.534 2.544 (+0.010)
We comprehensively analyze our model by answering three key questions on the 470M scale: (1) Are the proposed modules effective? (2) How efficient of TrinityDNA for long-sequence DNA computing? (3) Why do we use ETS for DNA modeling?
(1) Effectiveness.
Table 1 shows that GFM consistently lowers perplexity by modeling the spatial ‘groove’ features of DNA sequences. Likewise, incorporating GRC (which captures reverse-complement patterns) yields a notable drop in PPL, reflecting the importance of complementary-strand information. Meanwhile, SMWA enables multi-scale context handling, trading off some computational overhead for competitive perplexity across local and longer-range dependencies. SMWA also brings some performance gain due to the low-density perplexity of the nucleotide sequence.
(2) Efficiency.
The left panel of Figure 6 contrasts token-throughput as we sweep both sequence length and micro-batch size. Across every setting, TrinityDNA sits clearly at the top, retaining more than 80% of its short-sequence throughput even at 64 k tokens. This robustness comes from its sliding multi-window attention and optimized fused kernels, which keep memory traffic almost flat as context grows. By contrast, DNABERT-2 exhibits the steepest decline: its throughput falls by nearly an order of magnitude once the sequence length exceeds 16 k, reflecting the quadratic cost of its vanilla soft-max attention. The remaining baselines lie between these two extremes, showing gradual slowdowns that track their respective attention or RNN bottlenecks.
(3) Training Strategies.
The right plot in Figure 6 shows a comprehensive ablation study separating the contributions of dataset size and evolutionary training strategy. The table shows perplexity scores across different model configurations. Key finding: Models initialized with weights from prokaryotic pre-training and then fine-tuned on combined data show better performance than models trained from scratch on the combined dataset. This validates both the importance of large, diverse datasets and our evolutionary training approach.
4.3 Downstream Tasks
4.3.1 GUE Benchmark
Table 2: GUE Benchmark performance comparison across various datasets and tasks.
@@ 4.3.1 GUE Benchmark
H3 0.731 ±0.015$\pm$ 0.015± 0.015 0.788 ±0.012$\pm$ 0.012± 0.012 0.783 ±0.014$\pm$ 0.014± 0.014 0.799 ±0.029$\pm$ 0.029± 0.029 0.779 ±0.037$\pm$ 0.037± 0.037 0.814 ±0.014$\pm$ 0.014± 0.014
H3K14ac 0.401 ±0.018$\pm$ 0.018± 0.018 0.562 ±0.015$\pm$ 0.015± 0.015 0.526 ±0.019$\pm$ 0.019± 0.019 0.541 ±0.212$\pm$ 0.212± 0.212 0.612 ±0.065$\pm$ 0.065± 0.065 0.694 ±0.016$\pm$ 0.016± 0.016
H3K36me3 0.473 ±0.017$\pm$ 0.017± 0.017 0.620 ±0.012$\pm$ 0.012± 0.012 0.569 ±0.015$\pm$ 0.015± 0.015 0.609 ±0.109$\pm$ 0.109± 0.109 0.613 ±0.041$\pm$ 0.041± 0.041 0.692 ±0.014$\pm$ 0.014± 0.014
H3K4me1 0.414 ±0.012$\pm$ 0.012± 0.012 0.553 ±0.011$\pm$ 0.011± 0.011 0.505 ±0.015$\pm$ 0.015± 0.015 0.488 ±0.102$\pm$ 0.102± 0.102 0.512 ±0.024$\pm$ 0.024± 0.024 0.611 ±0.015$\pm$ 0.015± 0.015
H3K4me2 0.323 ±0.014$\pm$ 0.014± 0.014 0.365 ±0.010$\pm$ 0.010± 0.010 0.311 ±0.013$\pm$ 0.013± 0.013 0.388 ±0.101$\pm$ 0.101± 0.101 0.455 ±0.095$\pm$ 0.095± 0.095 0.480 ±0.013$\pm$ 0.013± 0.013
H3K4me3 0.278 ±0.015$\pm$ 0.015± 0.015 0.403 ±0.011$\pm$ 0.011± 0.011 0.363 ±0.014$\pm$ 0.014± 0.014 0.440 ±0.202$\pm$ 0.202± 0.202 0.549 ±0.056$\pm$ 0.056± 0.056 0.522 ±0.015$\pm$ 0.015± 0.015
H3K79me3 0.612 ±0.013$\pm$ 0.013± 0.013 0.647 ±0.010$\pm$ 0.010± 0.010 0.674 ±0.012$\pm$ 0.012± 0.012 0.676 ±0.026$\pm$ 0.026± 0.026 0.672 ±0.048$\pm$ 0.048± 0.048 0.741 ±0.014$\pm$ 0.014± 0.014
H3K9ac 0.512 ±0.015$\pm$ 0.015± 0.015 0.560 ±0.013$\pm$ 0.013± 0.013 0.556 ±0.017$\pm$ 0.017± 0.017 0.604 ±0.048$\pm$ 0.048± 0.048 0.581 ±0.061$\pm$ 0.061± 0.061 0.662 ±0.014$\pm$ 0.014± 0.014
H4 0.793 ±0.012$\pm$ 0.012± 0.012 0.817 ±0.015$\pm$ 0.015± 0.015 0.807 ±0.011$\pm$ 0.011± 0.011 0.789 ±0.020$\pm$ 0.020± 0.020 0.763 ±0.044$\pm$ 0.044± 0.044 0.829 ±0.012$\pm$ 0.012± 0.012
H4ac 0.372 ±0.016$\pm$ 0.016± 0.016 0.491 ±0.018$\pm$ 0.018± 0.018 0.504 ±0.019$\pm$ 0.019± 0.019 0.525 ±0.240$\pm$ 0.240± 0.240 0.564 ±0.038$\pm$ 0.038± 0.038 0.632 ±0.013$\pm$ 0.013± 0.013
Human TF 0.642 ±0.012$\pm$ 0.012± 0.012 0.633 ±0.015$\pm$ 0.015± 0.015 0.701 ±0.020$\pm$ 0.020± 0.020 - - 0.714 ±0.009$\pm$ 0.009± 0.009
Mouse TF 0.564 ±0.018$\pm$ 0.018± 0.018 0.670 ±0.014$\pm$ 0.014± 0.014 0.680 ±0.015$\pm$ 0.015± 0.015 - - 0.786 ±0.012$\pm$ 0.012± 0.012
Promoter 0.768 ±0.015$\pm$ 0.015± 0.015 0.799 ±0.012$\pm$ 0.012± 0.012 0.774 ±0.019$\pm$ 0.019± 0.019 - - 0.803 ±0.012$\pm$ 0.012± 0.012
Splice Recon. 0.841 ±0.010$\pm$ 0.010± 0.010 0.894 ±0.014$\pm$ 0.014± 0.014 0.850 ±0.020$\pm$ 0.020± 0.020 - - 0.927 ±0.009$\pm$ 0.009± 0.009
Virus Covid 0.555 ±0.017$\pm$ 0.017± 0.017 0.730 ±0.012$\pm$ 0.012± 0.012 0.710 ±0.014$\pm$ 0.014± 0.014 - - 0.706 ±0.015$\pm$ 0.015± 0.015

### v2

@@ 1 Introduction
The rapid advancements in large-scale, long-sequence modeling, particularly in the realm of Natural Language Processing (NLP) (Team et al. 2023; Achiam et al. 2023), have radically transformed the way we approach complex data. Deep learning models, such as Transformers (Vaswani et al. 2017), have achieved unprecedented success in tasks that span from language translation to text generation, revolutionizing not just NLP but a variety of other fields. These models have proven their capability to capture intricate dependencies in data, providing solutions to challenges that were once considered insurmountable. With these breakthroughs in NLP, there has emerged an exciting opportunity to extend the power of sequence modeling to a completely different domain—genomics—where data shares some key similarities, such as its sequential nature.
Genomic data, particularly DNA sequences, consists of extraordinarily long strings of information that encode the fundamental building blocks of life. Unlike the highly dense and structured data typically encountered in NLP, genomic sequences are sparse in nature, containing vast stretches of repetition and variability (Liu et al. 2024a). Despite this, they hold a rich repository of biological information that is crucial for understanding gene functions, regulatory mechanisms, and cellular processes. The ability to model DNA sequences deeply could lead to breakthrough applications in personalized medicine, genetic engineering, and the overall understanding of biological systems. However, effectively capturing the dependencies within such long, sparse sequences remains a significant challenge.
Figure 1: Overview of TrinityDNA Model: (Left) The evolutionary training strategy of TrinityDNA, progressing from prokaryotic DNA to multi-species eukaryotic DNA, and its DNA-targeted long-sequence modeling approach addressing structural features such as bidirectional complementarity and major/minor grooves. (Right) Radar chart illustrating the state-of-the-art performance on the zero-shot performance of our models versus popular models such as EVO and Caduceus.
While the parallels between NLP and genomics are evident, directly applying traditional NLP models to genomic sequences proves difficult. The sparse, low-density nature of DNA sequences means that existing models often struggle to identify long-range dependencies and interpret the underlying biological structures(Mallet and Vert 2021; Zhou et al. 2021). Moreover, the lack of biologically informed features in current models limits their effectiveness in genomic contexts. Many models trained on single-species data perform poorly when generalized to other species or broader biological contexts. As a result, the impact of these models on genomic research has been somewhat restricted, and their applicability to real-world challenges remains limited.
@@ 2 Background
Deoxyribonucleic acid (DNA) is the hereditary material in most living organisms, consisting of long sequences of nucleotides. Each nucleotide comprises a phosphate group, a deoxyribose sugar, and one of four nitrogenous bases: adenine (A), thymine (T), cytosine (C), and guanine (G). A DNA sequence can be represented as a string $S=(s_{1},s_{2},\ldots,s_{N})$, where $s_{i}\in\{A,T,C,G\}$ and $N$ denotes the sequence length.
@@
The double helix structure of DNA features two distinct grooves: the minor groove and the major groove. These grooves arise from the asymmetric positioning of the phosphate backbone relative to the base pairs. (1) Major Groove: wider and deeper, the major groove provides greater accessibility for protein binding and molecular interactions. It generally covers five to seven nucleotides. (2) Minor Groove: narrower and shallower, the minor groove presents a different arrangement of hydrogen bond donors and acceptors. While less accessible than the major groove, its length is three to five nucleotides.
@@
DNA molecules consist of two complementary strands running in opposite directions, a feature known as antiparallel orientation. For a DNA sequence $S=(s_{1},s_{2},\ldots,s_{N})$, its reverse complement $S^{R}$:
@@
DNA Long-Sequence Modeling
@@
A prominent class of models for handling long-range dependencies is based on Structured State Space Models (SSMs) (Gu et al. 2022a, 2021, b; Gupta et al. 2022; Smith et al. 2022; Dao et al. 2022). These models emerge from discretizing a continuous-time linear system:
@@
HyenaDNA (Poli et al. 2023) is an SSM variant, a decoder-style model capable of handling long genomic sequences (e.g., hundreds of thousands of tokens). It leverages the Hyena operator, which replaces traditional attention with a fast convolution-based mechanism. Concretely, HyenaDNA blocks compute a Toeplitz convolution filter on projected input tokens, allowing processing of very large contexts (in $\mathcal{O}(L\log L)$ time) without the quadratic cost typically associated with attention.
@@
MambaDNA (Schiff et al. 2024)—also referred to as Caduceus—builds on the selective SSM approach of Mamba and incorporates reverse-complement symmetry, a core property of DNA sequences. In standard Mamba, the module processes sequences in a single direction. MambaDNA extends this design in two ways: (1) BiMamba: Instead of only left-to-right processing, BiMamba applies the Mamba block twice. (2) RC Equivariance: MambaDNA explicitly enforces RC symmetry by taking a sequence and its RC as inputs to the same SSM-based module. See more details in Appendix E.
@@ 3 TrinityDNA: Sequence, Structure, and Strategy for DNA Modeling
Figure 2: Comparison of log influential scores log $|\partial yt/\partial xs|$ versus distance (ts)(t-s) on HG-38 (Nguyen et al. 2023).
Preliminaries
@@
While SSMs theoretically excel at handling long sequences, they inherently exhibit a locality bias (Wang et al. 2024). This issue is exacerbated in DNA sequence modeling, where dependencies across vast genomic regions must be captured to enable accurate biological interpretation. Specifically, in genomic data, long-range dependencies often span tens of thousands or even hundreds of thousands of base pairs. Existing models, such as SSMs, typically focus on local dependencies due to computational challenges in processing long sequences. Consequently, our empirical results in Figure 2 demonstrate that the SSM-based model (Caduceus) lost their focus as the sequence length increased, while the full-attention-based model (DNABERT2) suffered from heavy computation. These limitations hinder modern DNA foundation models’ ability to fully capture the inherent complexities of DNA sequences and their interconnections across genomic regions.
@@
As sequences grow in training, full self-attention “flattens out”: attention scores converge toward a uniform, high-entropy distribution in Figure 3, so every token is weighted almost equally, and useful signals vanish. This oversmoothing affects low-density data—images, DNA, and the hardest— where informative tokens are sparse and far apart. In shorter windows, the same model still shows specialized heads, e.g., retrieval vs. induction (Xiao et al. 2024), but that diversity collapses at the kilobase scale. These findings motivate a multi-window, multi-head scheme that combines narrow local windows with broader global ones to reduce entropy and maintain head specialization. Detailed settings refers to the Appendix.
@@
Groove Fusion Module
@@
Sliding Multi-Window Attention
To mitigate locality bias and attention oversmoothing, we revisit the design of Multi-Head Attention (MHA), focusing on varying attention window sizes across heads. In traditional MHA, each head typically attends to the entire sequence using a fixed attention window, which can limit the model’s ability to capture dependencies at different scales. To address this, we introduce a multi-scale attention grouping strategy named SMWA. In this approach, we assign different attention heads to capture dependencies at different scales in the DNA sequence, enabling the model to specialize in local or global dependencies, depending on the feature scale. Formally, we define the window sizes for each attention head $h$ as $L_{h}\in\mathbb{N}$, with $L_{h}$ representing the length of the attention window for head $h$. For head $h$, the attention mechanism is computed using a sliding window of size $L_{h}$ over the sequence, allowing the model to focus on scales:
$\text{Attn}_{h}(S_{i})=$ $\text{Softmax}\left(\frac{Q_{h}(i)K_{h}(i+[-L_{h},L_{h}])^{T}}{\sqrt{d_{k}}}\right)$
$V_{h}(i+[-L_{h},L_{h}])$
where $Q_{h}\in\mathbb{R}^{N\times d_{k}}$ is the query matrix for head $h$, $K_{h}\in\mathbb{R}^{N\times d_{k}}$ is the key matrix, $V_{h}\in\mathbb{R}^{N\times d_{v}}$ is the value matrix, $N$ is the sequence length, and $d_{k}$ and $d_{v}$ are the dimensionality of the key and value vectors, respectively. $i$ is the index of the sequence, and $L_{h},L_{h}]$ represents the range of indices for the sliding window around $i$. This enables each head to specialize in attending to either short-range or long-range dependencies by adjusting $L_{h}$. The final output of the multi-head attention layer is the concatenation of all heads’ outputs, followed by a linear transformation:
@@
Gated Reverse Complement
To leverage the reverse complementarity inherent in DNA sequences, we introduce a novel Gated Reverse Complement (GRC) mechanism. This mechanism employs a shared Transformer module to process both the forward and reverse complement sequences in parallel. The reverse complement of a DNA sequence $S=(s_{1},s_{2},\dots,s_{N})$ is defined as $S^{R}=(s_{N}^{C},s_{N-1}^{C},\dots,s_{1}^{C})$, where $s_{i}^{C}$ denotes the complementary base of $s_{i}$ (with base-pairing rules $A\leftrightarrow T$, $C\leftrightarrow G$). The GRC mechanism works by feeding both the forward and reverse complement sequences into a shared SMWA-equipped Transformer network $f_{\theta}$, where the outputs are gated using a linear gating mechanism to combine the two representations effectively. The final layer is:
@@
where $W_{G}$ are learned weights, $\sigma$ represents the identity function, and the Flip operator means to reverse the sequence in the original order. This allows the model to learn the representations of both sides simultaneously.
Evolutionary Training Strategy
The Evolutionary Training Strategy (ETS) approach leverages a two-stage, evolution-inspired training strategy to progressively address the varying complexities of genomic data. In the first stage, the model is trained on prokaryotic genomes, which are relatively straightforward in terms of their regulatory architectures (Xu and Jin 2006; Nguyen et al. 2024). Through this foundational phase, the model captures essential DNA sequence motifs and organizational patterns. Subsequently, the second stage introduces eukaryotic genomes, known for their intron-exon structures and gene lengths that can span tens of kilobases (Dalla-Torre et al. 2024). Alongside this transition, the model’s context window is enlarged from 8K to 100K base pairs, accommodating multiple co-expressed genes and regulatory elements.
@@ 4 Experiments
We adopt masked language modeling (MLM) (Devlin et al. 2018) and character-level tokenization for all our DNA models. Following our Evolutionary training strategy, we conduct pre-training in two phases: Stage 1 (Short-Sequence Pre-training): We use prokaryotic (bacterial and archaeal) DNA data from the OpenGenome dataset introduced in (Nguyen et al. 2024), training on sequences of length 8k to learn fundamental nucleic acid patterns in shorter contexts. Stage 2 (Multi-species Post-training): We then continue pre-training on a multi-species collection, the Multispecies dataset, presented in (Dalla-Torre et al. 2024). The sequences in this stage can be as long as 100k, spanning archaebacteria, fungi, vertebrate genomes, and more. This step exposes the model to a rich spectrum of evolutionary signals, enabling it to handle much longer contexts and to capture the diverse structural intricacies of eukaryotic DNA. Hence, we propose two main models, TrinityMicroDNA and TrinityDNA, each with 1 billion (1B) parameters. The former is trained solely on prokaryotic data, while the latter builds upon the former by post-training on eukaryotic data. By transitioning from bacterial genomes to longer eukaryotic genomes, our method achieves broad coverage of genomic features across scales. More details in Appendix A.
@@
Components W/O W W/O W
GRC 2.731 2.599 (-0.132) - -
GFM 2.599 2.534 (-0.065) - -
SMWA 2.534 2.544 (+0.010) 64.5 44.5 (-31%)
Table 1: Effect of components on pre-training perplexity (left side) and computational cost (left side, TFLOPs).
@@
H3 0.731 $\pm 0.015$ 0.788 $\pm 0.012$ 0.783 $\pm 0.014$ 0.799 $\pm 0.029$ 0.779 $\pm 0.037$ 0.814 $\pm 0.014$
H3K14ac 0.401 $\pm 0.018$ 0.562 $\pm 0.015$ 0.526 $\pm 0.019$ 0.541 $\pm 0.212$ 0.612 $\pm 0.065$ 0.694 $\pm 0.016$
H3K36me3 0.473 $\pm 0.017$ 0.620 $\pm 0.012$ 0.569 $\pm 0.015$ 0.609 $\pm 0.109$ 0.613 $\pm 0.041$ 0.692 $\pm 0.014$
H3K4me1 0.414 $\pm 0.012$ 0.553 $\pm 0.011$ 0.505 $\pm 0.015$ 0.488 $\pm 0.102$ 0.512 $\pm 0.024$ 0.611 $\pm 0.015$
H3K4me2 0.323 $\pm 0.014$ 0.365 $\pm 0.010$ 0.311 $\pm 0.013$ 0.388 $\pm 0.101$ 0.455 $\pm 0.095$ 0.480 $\pm 0.013$
H3K4me3 0.278 $\pm 0.015$ 0.403 $\pm 0.011$ 0.363 $\pm 0.014$ 0.440 $\pm 0.202$ 0.549 $\pm 0.056$ 0.522 $\pm 0.015$
H3K79me3 0.612 $\pm 0.013$ 0.647 $\pm 0.010$ 0.674 $\pm 0.012$ 0.676 $\pm 0.026$ 0.672 $\pm 0.048$ 0.741 $\pm 0.014$
H3K9ac 0.512 $\pm 0.015$ 0.560 $\pm 0.013$ 0.556 $\pm 0.017$ 0.604 $\pm 0.048$ 0.581 $\pm 0.061$ 0.662 $\pm 0.014$
H4 0.793 $\pm 0.012$ 0.817 $\pm 0.015$ 0.807 $\pm 0.011$ 0.789 $\pm 0.020$ 0.763 $\pm 0.044$ 0.829 $\pm 0.012$
H4ac 0.372 $\pm 0.016$ 0.491 $\pm 0.018$ 0.504 $\pm 0.019$ 0.525 $\pm 0.240$ 0.564 $\pm 0.038$ 0.632 $\pm 0.013$
Human TF 0.642 $\pm 0.012$ 0.633 $\pm 0.015$ 0.701 $\pm 0.020$ - - 0.714 $\pm 0.009$
Mouse TF 0.564 $\pm 0.018$ 0.670 $\pm 0.014$ 0.680 $\pm 0.015$ - - 0.786 $\pm 0.012$
Promoter 0.768 $\pm 0.015$ 0.799 $\pm 0.012$ 0.774 $\pm 0.019$ - - 0.803 $\pm 0.012$
Splice Recon. 0.841 $\pm 0.010$ 0.894 $\pm 0.014$ 0.850 $\pm 0.020$ - - 0.927 $\pm 0.009$
Virus Covid 0.555 $\pm 0.017$ 0.730 $\pm 0.012$ 0.710 $\pm 0.014$ - - 0.706 $\pm 0.015$